OpenStudy (anonymous):

8x+4y=12 solve for y

OpenStudy (anonymous):

8x + 4y = 12 Simplify by dividing all items by the GCF which is 4 2x + y = 3 y = 3 - 2x

OpenStudy (anonymous):

do you want why in numberical or an expression?